solve each equation and check.
-2/9=5x/6-1/3

Answer by haileytucki(390)

-(2)/(9)=(5x)/(6)-(1)/(3)
Since x is on the right-hand side of the equation, switch the sides so it is on the left-hand side of the equation.
(5x)/(6)-(1)/(3)=-(2)/(9)
Since -(1)/(3) does not contain the variable to solve for, move it to the right-hand side of the equation by adding (1)/(3) to both sides.
(5x)/(6)=(1)/(3)-(2)/(9)
Simplify the right-hand side of the equation.
(5x)/(6)=(1)/(9)
Multiply each term in the equation by 6.
(5x)/(6)*6=(1)/(9)*6
Simplify the left-hand side of the equation by canceling the common factors.
5x=(1)/(9)*6
Simplify the right-hand side of the equation by simplifying each term.
5x=(2)/(3)
Divide each term in the equation by 5.
(5x)/(5)=(2)/(3)*(1)/(5)
Simplify the left-hand side of the equation by canceling the common factors.
x=(2)/(3)*(1)/(5)
Simplify the right-hand side of the equation by simplifying each term.
x=(2)/(15)